OpenStudy (anonymous):

If 6 x 3y +6 = 89cm2 what is y

OpenStudy (anonymous):

18y=89-6=83 cm^2 y=83/18 cm^2

OpenStudy (mathmale):

Is that " x " in your expression the letter " x " or are you trying to indicate multiplication?

OpenStudy (mathmale):

Convention is that " x " is the letter x, and the multiplication operation is denoted by " * ".